# Data-Constrained Synthesis of Training Data for De-Identification

### v2

@@ 1 Introduction
Many useful applications in NLP involve domains where data are sensitive. These privacy risks, and the accompanying limits to sharing data, have traditionally been solved by de-identification. This process involves finding parts of the text that can be used to identify an individual. Such information is typically referred to as personally identifiable information (PII). After locating PII, the passages need to be processed to remove or obscure the PII. Traditionally, this time-consuming work has been done manually. Automatic de-identification Meystre et al. (2010) is a machine-driven approach that typically relies on named entity recognition (NER) to detect PII that needs to be removed.
Unfortunately, the PII datasets that exist to assist in privacy preservation are themselves sensitive, and can typically not be shared. This circularity, together with the increasing generative capabilities of large language models (LLMs), has led to a growing interest in overcoming data limitations by eschewing the use of real data altogether. Instead, one can use generated synthetic corpora.
Previous studies have mainly been concerned with evaluating the privacy of the synthetic text Yue et al. (2023); Miranda et al. (2024) or with creating the strongest-performing model possible using synthetic data Libbi et al. (2021); Hiebel et al. (2023); Liu et al. (2025). Our study instead examines how synthetic data can be produced under constrained resources. This understudied problem is common in clinical institutions that lack resources both in terms of data and computational hardware.
We carry out a systematic evaluation of key factors impacting the utility of LLM-generated synthetic data as training data for downstream tasks. Specifically, we study synthetic NER data for PII detection - an important task in the privacy-sensitive healthcare domain. Synthetic clinical data are generated using domain-adapted LLMs and machine-annotated using fine-tuned NER models. The evaluations focus primarily on the utility of synthetic data for training NER models.
@@ 2 Related Research
The second main current in the literature explores how to optimize synthesis to create the best possible synthetic corpora. These papers synthesize data using locally domain-adapted LLMs Ive et al. (2020); Hiebel et al. (2023), or using instruction-tuned models Kiefer (2024); Liu et al. (2025). They show that synthesis of high-utility data is possible. However, fewer papers systematically examine the conditions required for success.
@@ 3 Data and Methods
Figure 1: This figure illustrates the steps for creating the synthetic corpora. A gold standard corpus is used to train a NER model and to domain-adapt a general-domain LLM to produce synthetic clinical data. The LLM is used to generate synthetic data, and the NER model is used to add machine annotations. Later, these corpora are used to train synthetic NER models that are evaluated on gold standard test data.
@@ 3.1 Generative Models
Both models are used autoregressively, without instruction tuning. The hypothetical – but often occurring – scenario motivating the study design is where researchers have access to a small and sensitive NER dataset that cannot be shared outside of their organization. Zero-shot synthesis is an alternative strategy, but we leave to future research to evaluate if this approach can yield clinical texts that are sufficiently similar to the real data. Instead, we perform different degrees of domain-adaptive fine-tuning to train the LLMs to produce such texts.
@@ 3.2 NER Datasets
The Stockholm EPR PHI Pseudo Corpus (SEPR PHI) is a Swedish dataset from five different healthcare units consisting of 100 patient records split into 21,553 sentences. The corpus spans 282,766 tokens where 6,755 are manually annotated for nine different PII classes Velupillai et al. (2009). The dataset has then been pseudonymized, meaning that all the annotated entities have been replaced with realistic pseudonyms Dalianis (2019).
@@
The second dataset is MEDDOCAN – a Spanish dataset consisting of 1,000 medical texts. These are based on clinical cases augmented with PII from auxiliary sources Marimon et al. (2019). The texts were then manually annotated for 19 different PII. Out of 504,569 tokens, 41,859 are tagged as PII.
Each dataset was split into three subsets: one for training, another for validation, and a third held-out subset for testing. The training sets are used to domain-adapt the generative models and to train the machine-annotating NER models. The purpose of the validation subsets is two-fold. First, they are used to monitor the training processes when fine-tuning the models for synthesis and NER. Once the models for synthesis have been trained, the validation data also serve as starting points when creating the synthetic corpora. Finally, the quality of the NER models trained using the synthetic corpora is evaluated using the held-out test sets, as these are not in any way part of the training or synthesis processes.
@@ 3.3 Encoder Models
We use SweDeClin-BERT Vakili et al. (2022) for Swedish data, as it has previously shown strong performance on the SEPR PHI corpus. This BERT-style model is based on the general-domain KB-BERT model Devlin et al. (2019); Malmsten et al. (2020) and adapted to the clinical domain through continued pre-training on the Swedish Health Bank corpus Dalianis et al. (2015). It consists of 125 million parameters.
@@
These models are used for two purposes. First, they are fine-tuned using the gold standard datasets. These gold models are used to provide machine annotations to the synthesized corpora. They are also used as baselines for the later, synthetic NER models. The synthetic models are also initialized from the pre-trained encoder models, but are fine-tuned on the synthetic, machine-annotated corpora. The gold and synthetic NER models are then evaluated and compared to each other to measure the performance implications of using synthetic data.
@@ 3.4 Synthesizing and Evaluating the Corpora
The generative models were domain-adapted by fine-tuning them for causal language modeling using QLoRA Dettmers et al. (2023) as implemented in the Axolotl framework [^1] https://github.com/axolotl-ai-cloud/axolotl with $r=8$ and $\alpha=32$. A comprehensive specification of all hyperparameters is available in Appendix A. GPT-SW3 was domain adapted using the Swedish SEPR PHI corpus and FLOR using the Spanish MEDDOCAN corpus. Domain adaptation was carried out with varying amounts of data to determine the impact on the utility of the synthetic data.
@@
Synthesis was done using the vLLM library Kwon et al. (2023). We use nucleus sampling Holtzman et al. (2020) with $p=0.95$ and the minimum token length is set to 10. The maximum token length is set to the length of the longest document in the validation set, or at least 50. The temperature was set to $t=1.0$ after preliminary experiments showed that varying $0.8\leq t\leq 1.2$ had very little impact on the results.
@@
The whole process, from domain adaptation to training and evaluating the synthetic NER models, was done through five-fold cross-validation. The utility of the synthetic corpora was measured using token-level F1 score. These evaluations rely on each fold’s held-out gold-standard test data.
@@ 4 Experiments
The main contribution of this paper is a systematic investigation into how the different steps of synthetic corpora creation respond to data constraints. This section describes these experiments and their results. All experiments, except for those in Section 4.5, take advantage of the unbounded nature of synthetic corpora and allow them to be four times larger than the gold standard datasets. The performance of the NER models is summarized using token-level F1 scores tested on gold-standard data. All averages and standard deviations are calculated based on five-fold cross-validation.
@@ 4.1 Constraining the Total Amount of Data
Table 1: Between 5% and 95% of the training data in each fold was used to domain-adapt the generative LLMs and the machine-annotating encoder models. In this table, the F1 scores are listed for both the NER models trained on gold standard data and the synthetic data, as well as the difference. The F1 are the average scores over all 5 folds and the standard deviation is listed as well.
@@
Within each fold, this is scaled to between 5% and 95% of the training data in the fold. Five different amounts are used: 5%, 10%, 25%, 50%, and 95%. These subsets correspond to the Sensitive NER corpus in Figure 1. The final 5% are used for validation and for creating prompts for synthesis.
Synthetic corpora have an advantage over real corpora: their size is only constrained by the amount of compute available for generation. As explained in Section 3.4, this feature is incorporated by letting the synthetic data be four times larger than the original datasets. In later experiments, we analyze to what extent this advantage helps.
Table 1 lists the average F1 scores and their standard deviation for each tested configuration. Unsurprisingly, the performance of the gold models – the models trained on the real data – increases as more and more data are available for training. The MEDDOCAN models are more resilient to shrinking the amount of training data, but both cases show clear improvements as more data are available. The F1 scores of the models trained on synthetic corpora follow a similar pattern. Increasing the data allows for more data to domain-adapt the generative model and for training a better machine annotator.
@@ 4.2 Scaling the Data for Domain Adaptation
The experiments in the previous section show that it is indeed feasible to create well-performing NER models trained on synthetic data using our method. On the other hand, the results depend on the amount of data used for domain-adapting the synthesizing generative LLM and for fine-tuning the machine annotating encoder model. In the previous experiment, the data were fixed for both purposes.
@@ 4.3 Varying the Data for Machine Annotation
Table 3: The amount of data used to create the machine annotator (m.a.) varied between 5% and 95% of the training data in each fold. This table compares the downstream F1 scores of the synthetic and gold-standard NER models.
@@
In contrast to the previous experiments, these experiments use 95% of the data for domain adaptation of the generative model that produces the synthetic text. This corpus is still allowed to be four times larger than the original training corpus. The data used to create the machine annotating NER model is varied between 5% and 95%.
Models were trained and evaluated using 5-fold cross-validation and the resulting F1 scores are listed in Table 3. The average F1 scores adhere closely to those of the gold standard model that is trained on real data. This is especially clear when contrasting the scores with what was shown in Table 2. This strongly suggests that the performance of the synthetic models is mainly explained by the amount of data available when creating the machine annotators.
@@ 4.4 Using Smaller Generative Models
Gold 0.926 $\pm$ 0.005 0.973 $\pm$ 0.004
Table 4: GPT-SW3 and FLOR are available in smaller versions than those used in the other experiments. This table compares the downstream F1 scores obtained using the smaller and larger versions for domain adaptation.
@@ 4.6 Diversity of the Generated Data
Table 6: The lexical diversity, average length, and average number of annotated labels per document was calculated for the synthetic corpora and for the gold corpora. The synthetic corpora were machine annotated using models trained on 95% of the data. The domain adaptation (d.a.) varied between using 95% of the data, to using none.
@@
The lexical diversity of the synthetic data is fairly consistent, regardless of the amount of data used for domain adaptation. It is also consistently lower than in the gold corpora. This is likely due to the temperature being fixed across the configurations. As explained in Section 3.4, varying the temperature had a negligible impact on the downstream performance of the synthetic models. However, it is likely that the lexical diversity of the corpora would increase for higher temperatures.
@@ 4.7 Estimating Privacy
Creating a synthetic variant of a sensitive dataset only protects the original data if the synthetic and sensitive datasets are sufficiently different. A common proxy for measuring these risks is to study the n-grams of the original and synthetic corpora Ive et al. (2020); Hiebel et al. (2023). We calculated the n-gram recall of each generated dataset and the training data from which it is derived. This metric measures the proportion of unique n-grams in a reference document that is shared with a candidate document [^2] It is similar to ROUGE Lin (2004) but is used on a corpus level rather than for comparing individual documents.. In this experiment, the reference documents are the real sensitive corpora used for domain adaptation and the synthetic corpora are the candidates. In other words, given a real corpus with a set of n-grams $R$ and a synthetic corpus with n-grams $S$:
@@
Since the data in this study are tagged for PII, a PII-sensitive n-gram recall is also used to estimate the degree of leakage of potentially sensitive information. Instead of considering all n-grams $R$ in the reference document, as in Equation 1, this metric only considers the n-grams $R^{*}\subseteq R$ that overlap with sensitive entities in the gold standard corpus.
@@
N-gram recall values were calculated for $n=\{3,5,10\}$. This was done for each of the 5 folds in the previous experiments. The n-grams were created by concatenating $n$ tokens from the tokenizers of the domain-adapted generative models. The values for 5-grams are summarized in Table 7 and the other values in Tables 10 and 11 in Appendix B.
@@ 5 Discussion
An example of an application of our results is cross-institutional validation of NLP systems. In sensitive domains such as the clinical domain, researchers are often barred from sharing their data and trained models due to privacy concerns. A common is where research group (A) has created a system that works very well on their in-house data. Due to privacy regulations, another group (B) cannot share their data with (A), and this makes it difficult for (A) and (B) to validate if the system generalizes. An example of an attempt to work under these constraints is described by Bridal et al. (2022). They were able to make limited claims of generalization but where limited due to the restrictions on sharing data. The method for synthesis explored in our paper would allow these research groups to share synthetic versions of their datasets or models, as these are much less sensitive than artifacts based on real data.

### v3

@@ 1 Introduction
Many useful applications in NLP involve domains where the data are sensitive. These privacy risks, and the accompanying limits to sharing data, have traditionally been solved through de-identification. This process involves finding parts of the text that can be used to identify an individual. Such information is typically referred to as personally identifiable information (PII). After locating PII, the passages need to be processed to remove or obscure the PII. Traditionally, this time-consuming work has been done manually. Automatic de-identification Meystre et al. (2010) is a machine-driven approach that typically relies on named entity recognition (NER) to detect PII that needs to be removed.
Unfortunately, the PII datasets that exist to assist in privacy preservation are themselves sensitive and can usually not be shared. This circularity, together with the increasing generative capabilities of large language models (LLMs), has led to a growing interest in overcoming data limitations by eschewing the use of real data altogether. Instead, one can use generated synthetic corpora.
Previous studies have mainly been concerned with evaluating the privacy of the synthetic text Yue et al. (2023); Miranda et al. (2024) or with creating the strongest-performing model possible using synthetic data Libbi et al. (2021); Hiebel et al. (2023); Liu et al. (2025). Our study instead examines how synthetic data can be produced under constrained resources. This understudied problem is common in clinical institutions that lack resources, both in terms of data and computational hardware.
We carry out a systematic evaluation of key factors impacting the utility of LLM-generated synthetic data as training data for downstream tasks. Specifically, we study synthetic NER data for PII detection – an important task in the privacy-sensitive healthcare domain. Synthetic clinical data are generated using domain-adapted LLMs and machine-annotated using fine-tuned NER models. The evaluations focus primarily on the utility of synthetic data for training NER models.
@@ 2 Related Research
The second main current in the literature explores how to optimize synthesis to create the best possible synthetic corpora. These papers synthesize data using locally domain-adapted LLMs Ive et al. (2020); Hiebel et al. (2023), or using instruction-tuned models Kiefer (2024); Liu et al. (2025). They show that high-utility data synthesis is possible. However, fewer papers systematically examine the conditions required for success.
@@ 3 Data and Methods
Figure 1: This figure illustrates the steps for creating the synthetic corpora. A gold standard corpus is used to train a NER model and to domain-adapt a general-domain LLM to produce synthetic clinical data. The LLM is used to generate synthetic data, and the NER model is used to add machine annotations. Later, these corpora are used to train synthetic NER models that are evaluated on gold-standard test data.
@@ 3.1 Generative Models
Both models are used autoregressively, without instruction tuning. The hypothetical – but often occurring – scenario motivating the study’s design is where researchers have access to a small and sensitive NER dataset that cannot be shared outside of their organization. Zero-shot synthesis is an alternative strategy, but we leave it to future research to evaluate if this approach can yield clinical texts that are sufficiently similar to the real data. Instead, we perform different degrees of domain-adaptive fine-tuning to train the LLMs to produce such texts.
@@ 3.2 NER Datasets
The Stockholm EPR PHI [^1] Protected health information (PHI) is a term defined by the HIPAA regulation (HHS, 1996). The PHI term covers a subset of the types of information that constitute PII. Pseudo Corpus (SEPR PHI) is a Swedish dataset from five different healthcare units consisting of 100 patient records split into 21,553 sentences. The corpus spans 282,766 tokens, where 6,755 are manually annotated for nine different PII classes Velupillai et al. (2009). The dataset has then been pseudonymized, meaning that all the annotated entities have been replaced with realistic pseudonyms Dalianis (2019).
@@
The second dataset is MEDDOCAN – a Spanish dataset consisting of 1,000 medical texts. These are based on clinical cases augmented with PII from auxiliary sources Marimon et al. (2019). The texts were then manually annotated for 19 different PII. Out of 504,569 tokens, 41,859 are tagged as PII. The documents were originally divided into train, test, and development sets. In this paper, these three subsets have been merged before resampling them into new subsets for five-fold cross-validation.
Each dataset was split into three subsets: one for training, another for validation, and a third held-out subset for testing. The training sets are used to domain-adapt the generative models and to train the machine-annotating NER models. The purpose of the validation subsets is twofold. First, they are used to monitor the training processes when fine-tuning the models for synthesis and NER. Once the models for synthesis have been trained, the validation data also serve as starting points when creating the synthetic corpora. Finally, the quality of the NER models trained using the synthetic corpora is evaluated using the held-out test sets, as these are not in any way part of the training or synthesis processes.
@@ 3.3 Encoder Models
We use SweDeClin-BERT Vakili et al. (2022) for Swedish data as it has previously shown strong performance on the SEPR PHI corpus. This BERT-style model is based on the general-domain KB-BERT model Devlin et al. (2019); Malmsten et al. (2020) and adapted to the clinical domain through continued pre-training on the Swedish Health Bank corpus Dalianis et al. (2015). It consists of 125 million parameters.
@@
These models are used for two purposes. First, they are fine-tuned using the gold standard datasets. These gold models are used to add machine annotations to the synthesized corpora. They are also used as baselines for the later, synthetic NER models. The synthetic models are also initialized from the pre-trained encoder models, but are fine-tuned using the synthetic, machine-annotated corpora. The gold and synthetic NER models are then evaluated and compared to each other in order to measure the performance implications of using synthetic data.
@@ 3.4 Synthesizing and Evaluating the Corpora
The generative models were domain-adapted by fine-tuning them for autoregressive language modeling using QLoRA Dettmers et al. (2023) as implemented in the Axolotl framework [^2] https://github.com/axolotl-ai-cloud/axolotl with $r=8$ and $\alpha=32$. A comprehensive specification of all hyperparameters is available in Appendix A. GPT-SW3 was domain adapted using the Swedish SEPR PHI corpus and FLOR using the Spanish MEDDOCAN corpus. Domain adaptation was carried out with varying amounts of data to determine the impact on the utility of the synthetic data.
@@
Synthesis was done using the vLLM library Kwon et al. (2023). We use nucleus sampling Holtzman et al. (2020) with $p=0.95$, and the minimum token length is set to 10. The maximum token length is set to the length of the longest document in the validation set, or at least 50. The temperature was set to $t=1.0$ after preliminary experiments showed that varying $0.8\leq t\leq 1.2$ had very little impact on the results.
@@
The whole process, from domain adaptation to training and evaluating the synthetic NER models, was done through five-fold cross-validation. The utility of the synthetic corpora was measured using the token-level F1 score. These evaluations rely on each fold’s held-out gold-standard test data.
@@ 4 Experiments
The main contribution of this paper is a systematic investigation into how the different steps of synthetic corpora creation respond to data constraints. This section describes these experiments and their results. All experiments, except for those in Section 4.5, take advantage of the unbounded nature of synthetic corpora and allow them to be four times larger than the gold standard datasets. The performance of the NER models is summarized using token-level F1 scores tested on gold-standard data. All values are averages and standard deviations that are calculated based on the results from the five-fold cross-validation.
@@ 4.1 Constraining the Total Amount of Data
Table 1: Between 5% and 95% of the training data in each fold was used to domain-adapt the generative LLMs and the machine-annotating encoder models. In this table, the F1 scores are listed for both the NER models trained on gold standard data and the synthetic data, as well as the difference. The F1 scores are the average scores and their standard deviation over all five folds.
@@
Within each fold, this is scaled to between 5% and 95% of the training data in the fold. Four different amounts are used: 5%, 25%, 50%, and 95%. These subsets correspond to the Sensitive NER corpus in Figure 1. The final 5% are used for validation and for creating prompts for synthesis.
Synthetic corpora have an advantage over real corpora: their size is only constrained by the amount of computing power available for generation. As explained in Section 3.4, this feature is incorporated by letting the synthetic data be four times larger than the original datasets. In later experiments, we analyze the extent to which this advantage helps.
Table 1 lists the average F1 scores and their standard deviation for each tested configuration. Unsurprisingly, the performance of the gold models – the models trained on the real data – increases as more and more data are available for training. The MEDDOCAN models are more resilient to shrinking the amount of training data, but both cases show clear improvements as more data are available. The F1 scores of the models trained on synthetic corpora follow a similar pattern. Increasing the data allows for more data to be used to domain-adapt the generative model and for training a better machine annotator.
@@ 4.2 Scaling the Data for Domain Adaptation
The experiments in the previous section show that it is indeed feasible to create well-performing NER models trained on synthetic data using our method. On the other hand, the results depend on the amount of data used for domain-adapting the synthesizing generative LLM and for fine-tuning the machine-annotating encoder model. In the previous experiment, the data were fixed for both purposes.
@@ 4.3 Varying the Data for Machine Annotation
Table 3: The amount of data used to create the machine annotator (m.a.) varied between 5% and 95% of the training data in each fold. This table compares the downstream F1 scores of the synthetic and gold-standard NER models. The values are the average F1 scores and their standard deviation across all five folds.
@@
In contrast to the previous experiments, these experiments use 95% of the data for domain adaptation of the generative model that produces the synthetic text. This corpus is still allowed to be four times larger than the original training corpus. The data used to create the machine-annotating NER model is varied between 5% and 95%.
Models were trained and evaluated using five-fold cross-validation, and the resulting F1 scores are listed in Table 3. The average F1 scores adhere closely to those of the gold standard model that is trained on real data. This is especially clear when contrasting the scores with what was shown in Table 2. This strongly suggests that the performance of the synthetic models is mainly explained by the amount of data available when creating the machine annotators.
@@ 4.4 Using Smaller Generative Models
Gold 0.926 $\pm$ 0.005 0.978 $\pm$ 0.005
Table 4: GPT-SW3 and FLOR are available in smaller versions than those used in the other experiments. This table compares the average downstream F1 scores obtained using the smaller and larger versions for domain adaptation.
@@ 4.6 Diversity of the Generated Data
Table 6: The average lexical diversity, average length, and average number of annotated labels per document was calculated for the synthetic corpora and for the gold corpora. The synthetic corpora were machine annotated using models trained on 95% of the data. The domain adaptation (d.a.) varied between using 95% of the data, to using none.
@@
The lexical diversity of the synthetic data is fairly consistent, regardless of the amount of data used for domain adaptation. It is also consistently lower than in the gold corpora. This is likely due to the temperature being fixed across the configurations. As explained in Section 3.4, varying the temperature had a negligible impact on the downstream performance of the synthetic models. However, it is likely that the lexical diversity of the corpora would increase with higher temperatures.
@@ 4.7 Estimating Privacy
Creating a synthetic variant of a sensitive dataset only protects the original data if the synthetic and sensitive datasets are sufficiently different. A common proxy for measuring these risks is to study the n-grams of the original and synthetic corpora Ive et al. (2020); Hiebel et al. (2023). We calculated the n-gram recall of each generated dataset and the training data from which it is derived. This metric measures the proportion of unique n-grams in a reference document that is shared with a candidate document [^3] It is similar to ROUGE Lin (2004) but is used on a corpus level rather than for comparing individual documents.. In this experiment, the reference documents are the real sensitive corpora used for domain adaptation, and the synthetic corpora are the candidates. In other words, given a real corpus with a set of n-grams $R$ and a synthetic corpus with n-grams $S$:
@@
Since the data in this study are tagged for PII, a PII-sensitive n-gram recall is also used to estimate the degree of leakage of potentially sensitive information. Instead of considering all n-grams $R$ in the reference document, as in Equation 1, this metric only considers the n-grams $R^{*}\subseteq R$ that overlap with sensitive entities in the gold standard corpus. The relation between $R$, $R^{*}$, and $S$ is illustrated in Figure 2.
Figure 2: N-gram recall is calculated with all n-grams $R$ in the gold standard corpora, as well as with those in the subset $R^{*}\subseteq R$ that overlap with sensitive entities. The recall is calculated by counting how many of these are shared with $S$.
@@
N-gram recall values were calculated for $n=\{3,5,10\}$. This was done for each of the five folds in the previous experiments. The n-grams were created by concatenating $n$ tokens from the tokenizers of the domain-adapted generative models. The values for 5-grams are summarized in Table 7 and the other values in Tables 10 and 11 in Appendix B.
@@ 5 Discussion
An example of an application of our results is the cross-institutional validation of NLP systems. In sensitive domains such as the clinical domain, researchers are often barred from sharing their data and trained models due to privacy concerns. A common situation is when research group (A) has created a system that works very well on their in-house data. Due to privacy regulations, another group (B) cannot share their data with (A), and this makes it difficult for (A) and (B) to validate if the system generalizes. An example of an attempt to work under these constraints is described by Bridal et al. (2022). They were able to make limited claims of generalization, but were limited due to the restrictions on sharing data. The method for synthesis explored in our paper would allow these research groups to share synthetic versions of their datasets or models, as these are much less sensitive than artifacts based on real data.
Training models using synthetic corpora is safer than using real data, but is no panacea. On the other hand, no currently existing technique for privacy preservation is sufficient when used in isolation. For example, NER-based automatic de-identification covers only a subset of PII Pilán et al. (2022), and differentially private learning for NLP is difficult to implement properly Brown et al. (2022); Miranda et al. (2024) and is often inefficient when done so Igamberdiev et al. (2024). Synthetic data generation will likely be an important ingredient in many domains to overcome privacy issues.